# On Automating Video Game Testing by Planning and Learning

### v1

@@ Abstract
In this paper, we propose a method and workflow for automating the testing of certain video game aspects using automated planning and planning action model learning techniques. The basic idea is to generate detailed gameplay logs and apply action model learning to obtain a formal model in the planning domain description language (PDDL). The workflow enables efficient cooperation of game developers without any experience with PDDL or other formal systems and a person experienced with PDDL modeling but no game development skills. We describe the method and workflow in general and then demonstrate it on a concrete proof-of-concept example — a simple role-playing game provided as one of the tutorial projects in the popular game development engine Unity. This paper presents the first step towards minimizing or even eliminating the need for a modeling expert in the workflow, thus making automated planning accessible to a broader audience.
@@ Introduction
Planning is a less common skill game developers utilize but can be a powerful tool for game testing. One key aspect of that difficulty comes from the authorial burden of writing domain and problem descriptions in PDDL (Planning Domain Definition Language) (Aeronautiques et al. 1998). Good PDDL modeling is a rare skill that develops over time, so an efficient use of modelers by enabling them to be more effective is another key to adoption.
Our work addresses these critical issues to help game developers utilize planners for generating test scripts driven by PDDL through automation support. We accomplish this by taking the game logs, naturally produced in software development, providing easy-to-implement guidance to augment the logging for action model capture, and performing automated domain synthesis to generate PDDL models easily validated, verified, and adjusted by the game developer. These good PDDL examples also improve the developers’ understanding and authorial abilities in PDDL in the flow. Goal templates are then used to formulate domain problems of interest for testing. We demonstrate our approach and process on a simple RPG game in the ubiquitous Unity game engine (Unity Technologies 2023).
@@
This section briefly describes a few examples of what can be achieved when having a PDDL model of a video game in the testing context.
The most obvious usage is to generate a large amount of random test scenarios. This approach is beneficial if we can execute and evaluate test scenarios automatically. Next, we can search for dead-ends in the game, i.e., states that can be reached but are impossible to continue from or to win. In modern game design, such states are undesirable and annoying for the players. If the PDDL model is detailed enough, one can detect such dead ends by running planners from random reachable states. Last but not least, one can prove that undesired shortcuts are impossible through regular gameplay using optimal planners.
@@
Planning action model learning involves synthesizing a domain definition from logs with state sequences. Many algorithms have been published for this task (Wang 1996; Aineto, Celorrio, and Onaindia 2019; Zhuo et al. 2010; Yang, Wu, and Jiang 2007), a recent overview and new implementation of multiple techniques is provided in the MacQ project (Callanan et al. 2022). We use our own (publication pending) learning method in the evaluation, which most resembles the Observer algorithm (Wang 1996).
@@
Now, it is the modeler’s turn again, as they need to examine and evaluate the generated PDDL model manually. Due to the imprecise nature of AML algorithms, the generated model often needs minor corrections and adjustments. After such adjustments, it is possible to verify automatically whether the new PDDL model is still consistent with the logs. Note that the first model coming from the AML is always consistent with the logs. If the model is consistent with the logs, the modeler can either continue editing or decide that it is good enough and can be passed to the planning step. On the other hand, if the model is inconsistent with the logs, we need to either change the PDDL model or the logs. Changing the logs may be required because they are inconsistent due to bugs in the implementation of the logging process or because we are not logging the proper predicates. In the latter case, we may need to go back all the way to the beginning of the process and adjust the informal description, update the set of required predicates, and repeat all the follow-up steps (see Figure 1).
@@
It is a basic single-player RPG with a 2D environment viewed from above. The player controls a hero character using the arrow keys. The hero can talk to non-player characters (NPCs) on the map. Some NPCs can provide a “quest” (a task for the hero to complete) when spoken to. The quests all follow the pattern of: “collect items of type ”. After a quest is activated, the relevant items appear on the map. To pick up an item, the hero only needs to walk near it. When the required amount of items is collected, the player must visit the NPC who provided the quest to complete it. There are two quests in the demo: “collect 3 golden apples” and “collect 10 chickens.” Further data related to this example (such as logs, PDDL files, plans, and an execution video) is provided in the supplementary material of this submission.
@@
In this example the modeller decided to use a grid-based (i.e., via tiles) abstraction of locations111Using the grid abstraction turned out to be rather inconvenient. The reasons are that it is not very robust if the items are not aligned to the grid and its is hard to determine which tiles are connected. In a game with free movement such as our example it would be better to use a navigation-mesh-based abstraction (Shirzadehhajimahmood et al. 2021). and came up with following list of properties that should be logged:
@@
The developers then implemented logging and provided logs for learning. The modeller produced a PDDL domain and used a planner to a generate a plan to win the game. We used the FF planner (Hoffmann and Nebel 2001) which worked very well for our problem and found a plan in 0.01 seconds.
In order to test whether a generated plan (test script) is valid, we implemented automatic test script execution. In this case, it was very easy, since the game is controlled just by moving the hero. Quests are activated/completed automatically when their requirements are fulfilled and the hero comes near their NPC. Also, items are picked up automatically just by walking up to them. Therefore, to execute the test scripts we only need to execute the Move actions. We did that by simulating the pushing of the arrow buttons in the directions of the next goal tile until we are close enough to the center of that tile. Further data related to this example are provided in this submission’s supplementary material.
Challenges with Automated Logging and Learning
@@
Precisely, we must map all numeric variables, arrays, strings, and custom data structures to predicates. This step is because we cannot automatically decide what is necessary to describe all game-play rules fully. Interestingly, mapping all available variables is both too much detail and not necessarily sufficient at the same time. To completely model the behavior of a general computer program, we would also need to represent the “hidden” data like stack traces, the instruction pointer, the variables in the game engine, the operating system, and properties that are implied by the level design or interactions in the physics engine that are not represented in the code. For example, it is possible to jump from one location to another.
@@ Conclusion
In this paper we presented a workflow that applies research from the areas of automated planning and action model learning in order help with the automation of video game testing. We demonstrated it with a proof-of-concept project and discussed what challenges need to be addressed to further automate its steps.
@@
We plan to evaluate and refine the workflow on larger and more complex games. We also want to address and resolve the challenges related to further automating the synthesis of PDDL domains mentioned in the previous section.

### v2

# On Automating Video Game Regression Testing by Planning and Learning

@@ Abstract
In this paper, we propose a method and workflow for automating regression testing of certain video game aspects using automated planning and incremental action model learning techniques. The basic idea is to use detailed game logs and incremental action model learning techniques to maintain a formal model in the planning domain description language (PDDL) of the gameplay mechanics. The workflow enables efficient cooperation of game developers without any experience with PDDL or other formal systems and a person experienced with PDDL modeling but no game development skills. We describe the method and workflow in general and then demonstrate it on a concrete proof-of-concept example — a simple role-playing game provided as one of the tutorial projects in the popular game development engine Unity. This paper presents the first step towards minimizing or even eliminating the need for a modeling expert in the workflow, thus making automated planning accessible to a broader audience.
@@ Introduction
Planning is a less common skill game developers utilize but can be a powerful tool for game testing. One key aspect of that difficulty comes from the authorial burden of writing domain and problem descriptions in PDDL (Planning Domain Definition Language) (Ghallab et al. 1998). Good PDDL modeling is a rare skill that develops over time, so an efficient use of modelers by enabling them to be more effective is another key to adoption.
Our work addresses these critical issues to help game developers utilize planners for generating test scripts driven by PDDL through automation support. We accomplish this by taking the game logs, produced in software development, providing easy-to-implement guidance to augment the logging for action model capture, and performing automated domain synthesis to generate PDDL models easily validated, verified, and adjusted by the game developer. These good PDDL examples also improve the developers’ understanding and authorial abilities in PDDL in the flow. Goal templates are then used to formulate domain problems of interest for testing. We demonstrate our approach and process on a simple RPG game in the ubiquitous Unity game engine (Unity Technologies 2023).
@@
We will focus on regression testing during game development, i.e., running frequent tests to ensure that recent changes did not introduce unintended breaks and the game still performs as expected. This is a particularly repetitive and dull task for a human tester but crucial for efficient development. It is often the case, that longer the time difference between the introduction and discovery of the bug, the harder it is to fix.
What can be tested with Planning
The most obvious and important usage is to test whether the game can be completed, i.e., each intended objective can be achieved from a starting state of each scenario. Next, we can generate a large amount of random test scenarios. This approach is beneficial if we can execute and evaluate test scenarios automatically. Furthermore, we can search for dead-ends in the game, i.e., states that can be reached but are impossible to continue from or to win. In modern game design, such states are most undesirable and annoying for the players. If the PDDL model is detailed enough, one can detect such dead ends by running planners from random reachable states. Last but not least, one can prove that undesired shortcuts are impossible through regular gameplay using optimal planners.
Applicability
The classical planning based testing approach is most suitable for games with many discreet causal interactions such as role playing games (RPG), point-and-click adventures, strategy games, visual novels or puzzle games. It is less suitable for fast-paced action oriented games such as shooters or driving games. For such games, reinforcement learning based approaches, that have been studied in previous work, seem to better suitable (see the Related Work Section of this paper). Overall, we observe that the reinforcement learning and the planning based approaches can complement each other rather well.
@@
Planning action model learning (AML) involves synthesizing a domain definition from logs with state sequences. Many algorithms have been published for this task (Wang 1996; Aineto, Celorrio, and Onaindia 2019; Zhuo et al. 2010; Yang, Wu, and Jiang 2007; Juba, Le, and Stern 2021), a recent overview and new implementation of multiple techniques is provided in the MacQ project (Callanan et al. 2022). For the evaluation in this paper we use our own new action learning method (Balyo et al. 2024).
@@
Many games already produce logs, so why cannot we just use existing logs instead of modifying the code and creating new ones? In principle, we could do that. However, in our experience, the existing logs (if they are available at all) usually only contain method calls and error messages and do not record changes of state variables. Those kinds of logs unfortunately cannot be used to learn action models because the required information (state changes) is not present in them.
Now, it is the modeler’s turn again, as they need to examine and evaluate the generated PDDL model manually. Due to the imprecise nature of AML algorithms, the generated model often needs minor corrections and adjustments. After such adjustments, it is possible to verify automatically whether the new PDDL model is still consistent with the logs. Note that the first model coming from the AML is always consistent with the logs (this is a property of the used action model learning algorithm (Balyo et al. 2024)). If the model is consistent with the logs, the modeler can either continue editing or decide that it is good enough and can be passed to the planning step. On the other hand, if the model is inconsistent with the logs, we need to either change the PDDL model or the logs. Changing the logs may be required because they are inconsistent due to bugs in the implementation of the logging process or because we are not logging the proper predicates. In the latter case, we may need to go back all the way to the beginning of the process and adjust the informal description, update the set of required predicates, and repeat all the follow-up steps (see Figure 1).
@@
It is a basic single-player RPG with a 2D environment viewed from above. The player controls a hero character using the arrow keys. The hero can talk to non-player characters (NPCs) on the map. Some NPCs can provide a “quest” (a task for the hero to complete) when spoken to. The quests all follow the pattern of: “collect items of type ”. After a quest is activated, the relevant items appear on the map. To pick up an item, the hero only needs to walk near it. When the required amount of items is collected, the player must visit the NPC who provided the quest to complete it. There are two quests in the demo: “collect 3 golden apples” and “collect 10 chickens”.
@@
In this example the modeller decided to use a grid-based (i.e., via tiles) abstraction of locations111Using the grid abstraction turned out to be rather inconvenient. The reasons are that it is not very robust if the items are not aligned to the grid and its is hard to determine which tiles are connected. In a game with free movement such as our example it would be better to use a navigation-mesh-based abstraction (Shirzadehhajimahmood et al. 2021). and came up with the following list of properties that should be logged:
@@
To see the concrete predicate declarations corresponding to these properties see Figure 3.
The developers then implemented logging and provided logs for learning (see Figure 2). The modeller refined the PDDL domain synthesized from the logs (see Figures 3 and 4) and used a planner to a generate a plan to win the game (see Figure 6). We used the FF planner (Hoffmann and Nebel 2001) which worked very well for our problem and found a plan in 0.01 seconds.
In order to test whether a generated plan (test script) is valid222A test script is valid if it is executable in a bug-free implementation of the game. If a test script is valid, but not executable, then a bug in the game is indicated, we implemented automatic test script execution. In this case, it was very easy, since the game is controlled just by moving the hero. Quests are activated/completed automatically when their requirements are fulfilled and the hero comes near their NPC. Also, items are picked up automatically just by walking up to them. Therefore, to execute the test scripts we only need to execute the Move actions. We did that by simulating the pushing of the arrow buttons in the directions of the next goal tile until we are close enough to the center of that tile333A recording of the plan execution within the game is available here: https://www.youtube.com/watch?v=BASKvQAbG04.
Usage During Development
The process we described should be applied in the beginning of the development process, when the game is simple and only contains a few game mechanics like our RPG demo. As the game is developed and new features are added, the PDDL model should be developed together with the game. Using incremental action model learning can aid the developers with maintaining the PDDL model.
Challenges with Fully Automated Logging for Learning
@@
Precisely, we must map all numeric variables, arrays, strings, and custom data structures to predicates. This step is required because we cannot automatically decide what is necessary to describe all game-play rules fully. Interestingly, mapping all available variables is both too much detail and not necessarily sufficient at the same time. To completely model the behavior of a general computer program, we would also need to represent the “hidden” data like stack traces, the instruction pointer, the variables in the game engine, the operating system, and properties that are implied by the level design or interactions in the physics engine that are not represented in the code. For example, it is possible to jump from one location to another.
@@ Conclusion
In this paper we presented a workflow that applies research from the areas of automated planning and action model learning in order help with the automation of regression testing during video game development. We demonstrated it with a proof-of-concept project and discussed what challenges need to be addressed to further automate its steps.
@@
We plan to evaluate and refine the workflow on larger and more complex games. We are currently experimenting by applying it on a real-time strategy game. We also want to run proper user studies to quantify the usefulness of the proposed workflow compared to just creating the domains manually. Furthermore, we want to address and resolve the challenges related to further automating the synthesis of PDDL domains mentioned in the previous section.